lET a(Z1), b (Z2), c(Z3) , AND d(Z4) be the vertices of a trapezium in an argand plane. Let |z1-z2|-4 and |z3-z4|=10 and diangonals ac and bd intersect at p. It is given that arg(z4-z3/z3-z1)=pi/2 and arg(z3-z2/z4-z1)=pi/4.
 
Find the area of trapezium ABCD.
 
Ans>> 140/3

Grade:8

3 Answers

Jee
9 years ago
Is this question so tough to answer? Its been 5 days
its my doubt also...
Prashmitha
3 years ago
triangle abp and triangle cdp are similar,
   if ap=2x,bp=2y
then cp=5x,dp=5y
.: area of ABCD =(49xy)/2
tan alpha=2x/5y,tan beta=2y/2x also alpha+beta=45
xy=(10/21)x^2+y^2
also,ab^2=ap^2+bp^2
       x^2+y^2=4
        xy=40/21
.: area of trapezium=(49/2)(40/21)=140/3
 
Vikas TU
3 years ago
Dear student 
Question is not clear 
Please upload an image.
We will happy to help you.
Good luck 
Cheers